OpenStudy (anonymous):

HELP! Convergent, divergent, absolutely convergent, or conditionally convergent. Show all work and state what test you used to make your determination.

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty}(-1)^n \frac{ 1 }{ \sqrt[n]{n} }\]

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty}\frac{ n+\cos n }{ n^3+1 }\]